A sample of radioactive material has a half-life of 35 days. Calculate the fraction of the quantity that will remain after 105days
1
Expert's answer
2021-07-14T10:44:43-0400

105 days is equal to 3 half-lives ("\frac{t}{t_{1/2}}=\frac{105}{35}=3" ).

Assuming the initial quantity (N0) of the material to be 1, the fraction equals:

"N=N_o(\frac{1}{2})^{\frac{t}{t_{1/2}}}=(\frac{1}{2})^3=0.125"


Answer: 0.125